OpenStudy (anonymous):

Which property justifies the following equation? 7(6+0+5)=(6+5) A. distributive B. commutative C. associative D. identity ???

OpenStudy (mathstudent55):

You mean 7(6 + 0 + 5) = 7(6 + 5)? Did you leave the 7 out? 0 is identy for addition, a + 0 = a

OpenStudy (anonymous):

oops i did leave the 7 out sorry. :/ thanks though!